# On the Convergence Rates of Federated Q-Learning across Heterogeneous Environments

### v1

@@ Appendix F Proof of Theorem 2
$Q_{(r+1)E}^{k}=(A^{k})^{E}Q_{rE}^{k}+\left((A^{k})^{0}+\dots(A^{k})^{E-1}\right)\lambda R.$
@@
$\bar{Q}_{(r+1)E}=\bar{A}^{(E)}\bar{Q}_{rE}+\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)Q^{*}.$
@@
$\Delta_{(r+1)E}$ $=Q^{*}-\bar{Q}_{(r+1)E}$
$=\bar{A}^{(E)}\Delta_{rE}+\left(\left(I-\bar{A}^{(E)}\right)-\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)\right)Q^{*}.$ (28)
@@
$\frac{1}{k}\sum_{k=1}^{K}(P^{k})^{\ell}=\begin{cases}I,&\ell~{}\text{even},\\ \bar{P},&\ell~{}\text{odd}.\\ \end{cases}$
Applying the definition of $\bar{A}^{(\ell)}$ in equation (27) yields that
@@
Note that $0\leq\alpha\leq\beta\leq 1$ and $I-\bar{P}$ and $\bar{P}$ are orthogonal projection matrices satisfying $(I-\bar{P})\bar{P}=0$. The matrices for the second term of the error on the right-hand side of 28 reduce to
[…]
@@
[…]
@@
Applying equation (29) yields that
@@
It follows from equation (28) that the error evolves as
$\Delta_{(r+1)E}=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)\Delta_{rE}+\kappa_{E}(I-\bar{P})Q^{*},$
@@
$\Delta_{rE}$ $=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)^{r}\Delta_{0}+\sum_{\ell=0}^{r-1}\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)^{\ell}\kappa_{E}(I-\bar{P})Q^{*}$
$=\left(\alpha_{E}^{r}(I-\bar{P})+\beta_{E}^{r}\bar{P}\right)\Delta_{0}+\sum_{\ell=0}^{r-1}\left(\alpha_{E}^{\ell}(I-\bar{P})+\beta_{E}^{\ell}\bar{P}\right)\kappa_{E}(I-\bar{P})Q^{*}$
[…]
$=\left(\alpha_{E}^{r}(I-\bar{P})+\beta_{E}^{r}\bar{P}\right)\Delta_{0}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}(I-\bar{P})Q^{*}$
$=\left(\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}\right)(I-\bar{P})Q^{*}+\beta_{E}^{r}\bar{P}Q^{*},$
@@
$\left\|{\Delta_{rE}}\right\|_{\infty}\geq\frac{1}{\sqrt{2}}\left\|{\Delta_{rE}}\right\|_{2}\geq\frac{\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}}{\sqrt{2}}\cdot\max\left\{|\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{r}\right\}.$
Since $Q^{*}=(I-\gamma\bar{P})^{-1}R=(I-\bar{P})R+\frac{1}{1-\gamma}\bar{P}R$, we obtain that
$\|(I-\bar{P})Q^{*}\|_{2}=\|(I-\bar{P})R\|_{2},\qquad\|\bar{P}Q^{*}\|_{2}=\frac{1}{1-\gamma}\|\bar{P}R\|_{2}.$
Therefore, for the reward $R$ in general position, we have $\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}\geq c_{R}$ for some constant $c_{R}$ depending on the reward function. It remains to analyze the coefficients as functions of $\lambda$. To this end, we introduce the following lemma:
@@
    Note that $\nu_{1}<\nu_{2}$, $1-\nu_{1}=(1+\gamma)\lambda$, and $1-\nu_{2}=(1-\gamma)\lambda$. Then it follows from equation (30) that
@@
[…]
@@
We consider two regimes of the stepsize separated by $\lambda_{0}\triangleq\frac{\log r}{(1-\gamma)rE}<\frac{1}{1+\gamma}$, where the dominating error is due to the small stepsize and the environment heterogeneity, respectively:
@@
Since $\beta_{E}^{r}$ monotonically decreases as $\lambda$ increases,
$\beta_{E}^{r}=(1-(1-\gamma)\lambda)^{rE}\geq(1-(1-\gamma)\lambda_{0})^{rE}=\left(1-\frac{\log r}{rE}\right)^{rE}.$
Note that $\frac{\log r}{rE}\in(0,\frac{1}{2})$, applying the fact $\log(1-x)+x\geq-x^{2}$ for $x\in[0,\frac{1}{2}]$ yields that
$\log\left(1-\frac{\log r}{rE}\right)+\frac{\log r}{rE}\geq-\left(\frac{\log r}{rE}\right)^{2}\geq-\frac{1}{rE}.$
@@
$\beta_{E}^{r}\geq\left(1-\frac{\log r}{rE}\right)^{rE}\geq\frac{1}{er}.$
@@
$\left|\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}\right|\geq\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|-\left(1+\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\right)\alpha_{E}^{r}.$
For the first term, by the negativety and monotonicity in Lemma 5, it suffices to show the lower bound at $\lambda=\lambda_{0}$. Since $\lambda<\frac{1}{1+\gamma}$, then $\alpha_{E}=\frac{1}{2}\left((1-(1-\gamma)\lambda)^{E}+(1-(1+\gamma)\lambda)^{E}\right)$ decreases as $\lambda$ increases. For $r\geq\exp\left\{-W_{-1}\left(-\frac{1-\gamma}{2(1+\gamma)}\right)\right\},\text{ where }W_{-1}\text{ is the Lambert $W$ function}$, such that $(1+\gamma)\lambda_{0}\leq\frac{1}{2E}$, we apply the lower bound in Lemma 5 and obtain that
$\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\geq\frac{\lambda_{0}\gamma^{2}(E-1)}{4}\geq\frac{\frac{\log r}{(1-\gamma)rE}\gamma^{2}(E-1)}{4}\geq\frac{(E-1)}{4E}\gamma^{2}\frac{\log r}{(1-\gamma)r}$
@@
$\left(1+\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\right)\alpha_{E}^{r}\leq\frac{\nu_{2}^{rE}}{1-\gamma^{2}}=\frac{(1-(1-\gamma)\lambda)^{rE}}{1-\gamma^{2}}\leq\frac{1}{(1-\gamma^{2})r}$
@@
$\left|{\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}}\right|$ $\geq\left|{\frac{\kappa_{E}}{1-\alpha_{E}}}\right|-\left({1+\left|{\frac{\kappa_{E}}{1-\alpha_{E}}}\right|}\right)\alpha_{E}^{r}$
$\geq\frac{(E-1)}{4E}\gamma^{2}\frac{\log r}{(1-\gamma)r}-\frac{1}{(1-\gamma^{2})r}$
$=\frac{1}{(1-\gamma^{2})r}\left({(1+\gamma)\gamma^{2}\log(r)(E-1)/(4E)-1}\right)$
$=\frac{1}{(1-\gamma)r}\left({\frac{\gamma^{2}\log(r)(E-1)-4E/(1+\gamma)}{4E}}\right)$
Finally,
$\|{\Delta_{rE}}\|_{\infty}$ $\geq\frac{\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}}{\sqrt{2}}\cdot\max\left\{|\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{r}\right\}$
$\geq\frac{c_{R}}{\sqrt{2}}\max\left\{|\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{r}\right\}$
$\geq\frac{c_{R}}{\sqrt{2}}\max\left\{\frac{1}{(1-\gamma)r}\left(\frac{\gamma^{2}\log(r)(E-1)-4E/(1+\gamma)}{4E}\right),\frac{1}{er}\right\}$
$=\frac{c_{R}}{\sqrt{2}}\max\left\{\frac{E}{(1-\gamma)T}\left(\frac{\gamma^{2}\log(r)(E-1)-4E/(1+\gamma)}{4E}\right),\frac{E}{eT}\right\}$
We can choose $\log(r)\geq\frac{4E(\gamma+2)}{(1+\gamma)\gamma^{2}(E-1)},E\geq 2$ so that $\left(\frac{\gamma^{2}\log(r)(E-1)-4E/(1+\gamma)}{4E}\right)\geq 1$. Then the first term inside the max operator is bigger, and $c_{R}$ is also a numeric constant when $R$ is given. Therefore, the lower bound can be expressed as
$\|\Delta_{rE}\|_{\infty}\geq\frac{c_{R}}{\sqrt{2}}\frac{E}{(1-\gamma)T}\left(\frac{\gamma^{2}\log(r)(E-1)-4E/(1+\gamma)}{4E}\right)\geq\frac{c_{R}}{\sqrt{2}}\frac{E}{(1-\gamma)T}$
@@ Appendix G Additional experiments
(d) E=$\infty$
Figure 5: Homogeneous FQL with varying $E$.
G.2 Different target error levels.

### v2

@@ Appendix F Proof of Theorem 2
$Q_{(z+1)E}^{k}=(A^{k})^{E}Q_{zE}^{k}+\left((A^{k})^{0}+\dots(A^{k})^{E-1}\right)\lambda R.$
@@
$\bar{Q}_{(z+1)E}=\bar{A}^{(E)}\bar{Q}_{zE}+\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)Q^{*}.$
@@
$\Delta_{(z+1)E}$ $=Q^{*}-\bar{Q}_{(z+1)E}$
$=\bar{A}^{(E)}\Delta_{zE}+\left(\left(I-\bar{A}^{(E)}\right)-\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)\right)Q^{*}.$ (33)
@@
[…]
Applying the definition of $\bar{A}^{(\ell)}$ in (32) yields that
@@
Note that $0\leq\alpha\leq\beta\leq 1$ and $I-\bar{P}$ and $\bar{P}$ are orthogonal projection matrices satisfying $(I-\bar{P})\bar{P}=0$. The matrices for the second term of the error on the right-hand side of 33 reduce to
$\phantom{{}={}}\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)$
@@
$\phantom{{}={}}\left(I-\bar{A}^{(E)}\right)-\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)$
@@
Applying (34) yields that
@@
It follows from (33) that the error evolves as
$\Delta_{(z+1)E}=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)\Delta_{zE}+\kappa_{E}(I-\bar{P})Q^{*},$
@@
$\Delta_{zE}$ $=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)^{z}\Delta_{0}+\sum_{\ell=0}^{z-1}\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)^{\ell}\kappa_{E}(I-\bar{P})Q^{*}$
$=\left(\alpha_{E}^{z}(I-\bar{P})+\beta_{E}^{z}\bar{P}\right)\Delta_{0}+\sum_{\ell=0}^{z-1}\left(\alpha_{E}^{\ell}(I-\bar{P})+\beta_{E}^{\ell}\bar{P}\right)\kappa_{E}(I-\bar{P})Q^{*}$
 $\phantom{{}={}}\text{ since }\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)^{\ell}=\alpha_{E}^{\ell}(I-\bar{P})+\beta_{E}^{\ell}\bar{P},\forall\ell\in\mathbb{N}$
$=\left(\alpha_{E}^{z}(I-\bar{P})+\beta_{E}^{z}\bar{P}\right)\Delta_{0}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}(I-\bar{P})Q^{*}$
$=\left(\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}\right)(I-\bar{P})Q^{*}+\beta_{E}^{z}\bar{P}Q^{*},$
@@
$\left\|{\Delta_{zE}}\right\|_{\infty}\geq\frac{1}{\sqrt{2}}\left\|{\Delta_{zE}}\right\|_{2}\geq\frac{\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}}{\sqrt{2}}\cdot\max\left\{|\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{z}\right\}.$
Let $R=\begin{bmatrix}r_{1}\\ r_{2}\end{bmatrix}$, since $Q^{*}=(I-\gamma\bar{P})^{-1}R=(I-\bar{P})R+\frac{1}{1-\gamma}\bar{P}R$, we obtain that
$(I-\bar{P})Q^{*}=(I-\bar{P})R=\frac{1}{2}\begin{bmatrix}r_{1}-r_{2}\\ r_{2}-r_{1}\end{bmatrix},\qquad\bar{P}Q^{*}=\frac{1}{1-\gamma}\bar{P}R=\frac{1}{2(1-\gamma)}\begin{bmatrix}r_{1}+r_{2}\\ r_{1}+r_{2}\end{bmatrix}$.
$\|(I-\bar{P})Q^{*}\|_{2}=\frac{\sqrt{2}}{2}\left|r_{1}-r_{2}\right|,\qquad\|\bar{P}Q^{*}\|_{2}=\frac{\sqrt{2}}{2(1-\gamma)}\left|r_{1}+r_{2}\right|.$
When $r_{1}=r_{2},$ the error $\Delta_{zE}$ reduces to $\beta_{E}^{z}\bar{P}Q^{*}$, and $\left\|{\Delta_{zE}}\right\|_{\infty}=\frac{1}{2(1-\gamma)}\left|r_{1}+r_{2}\right|\left|\beta_{E}^{z}\right|$; otherwise, $\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}=\frac{\sqrt{2}}{2}\min\{\left|r_{1}-r_{2}\right|,\frac{1}{1-\gamma}\left|r_{1}+r_{2}\right|\}.$ It remains to analyze the coefficients as functions of $\lambda$. To this end, we introduce the following lemma:
@@
    Note that $\nu_{1}<\nu_{2}$, $1-\nu_{1}=(1+\gamma)\lambda$, and $1-\nu_{2}=(1-\gamma)\lambda$. Then it follows from (35) that
@@
    $\frac{\sum_{\ell=1}^{E-1}(\nu_{2}^{\ell}-\nu_{1}^{\ell})}{(1-\nu_{1}^{E})+(1-\nu_{2}^{E})}\geq\frac{\sum_{\ell=1}^{E-1}\int_{\nu_{1}}^{\nu_{2}}\ell\cdot x^{\ell-1}dx}{2\lambda E}\geq\frac{\sum_{\ell=1}^{E-1}\ell\frac{1}{2}(\nu_{2}-\nu_{1})}{2\lambda E}=\frac{1}{4}\gamma(E-1).$
@@
We consider two regimes of the stepsize separated by $\lambda_{0}\triangleq\frac{\log z}{(1-\gamma)zE}<\frac{1}{1+\gamma}$, where the dominating error is due to the small stepsize and the environment heterogeneity, respectively:
@@
Since $\beta_{E}^{z}$ monotonically decreases as $\lambda$ increases,
$\beta_{E}^{z}=(1-(1-\gamma)\lambda)^{zE}\geq(1-(1-\gamma)\lambda_{0})^{zE}=\left(1-\frac{\log z}{zE}\right)^{zE}.$
Note that $\frac{\log z}{zE}\in(0,\frac{1}{2})$, applying the fact $\log(1-x)+x\geq-x^{2}$ for $x\in[0,\frac{1}{2}]$ yields that
$\log\left(1-\frac{\log z}{zE}\right)+\frac{\log r}{zE}\geq-\left(\frac{\log z}{zE}\right)^{2}\geq-\frac{1}{zE}.$
@@
$\beta_{E}^{z}\geq\left(1-\frac{\log z}{zE}\right)^{zE}\geq\frac{1}{ez}.$
@@
$\left|\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}\right|\geq\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|-\left(1+\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\right)\alpha_{E}^{z}.$
For the first term, by the negativity and monotonicity in Lemma 5, it suffices to show the lower bound at $\lambda=\lambda_{0}$. Since $\lambda<\frac{1}{1+\gamma}$, then $\alpha_{E}=\frac{1}{2}\left((1-(1-\gamma)\lambda)^{E}+(1-(1+\gamma)\lambda)^{E}\right)$ decreases as $\lambda$ increases. For $z\geq\exp\left\{-W_{-1}\left(-\frac{1-\gamma}{2(1+\gamma)}\right)\right\},\text{ where }W_{-1}\text{ is the Lambert $W$ function}$, such that $(1+\gamma)\lambda_{0}\leq\frac{1}{2E}$, we apply the lower bound in Lemma 5 and obtain that
$\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\geq\frac{\lambda_{0}\gamma^{2}(E-1)}{4}\geq\frac{\frac{\log z}{(1-\gamma)zE}\gamma^{2}(E-1)}{4}\geq\frac{(E-1)}{4E}\gamma^{2}\frac{\log z}{(1-\gamma)z}.$
@@
$\left(1+\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\right)\alpha_{E}^{z}\leq\frac{\nu_{2}^{zE}}{1-\gamma^{2}}=\frac{(1-(1-\gamma)\lambda)^{zE}}{1-\gamma^{2}}\leq\frac{1}{(1-\gamma^{2})z}.$
@@
$\left|{\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}}\right|$ $\geq\left|{\frac{\kappa_{E}}{1-\alpha_{E}}}\right|-\left({1+\left|{\frac{\kappa_{E}}{1-\alpha_{E}}}\right|}\right)\alpha_{E}^{z}$
$\geq\frac{(E-1)}{4E}\gamma^{2}\frac{\log z}{(1-\gamma)z}-\frac{1}{(1-\gamma^{2})z}$
$=\frac{1}{(1-\gamma^{2})z}\left({(1+\gamma)\gamma^{2}\log(z)(E-1)/(4E)-1}\right)$
$=\frac{1}{(1-\gamma)z}\left({\frac{\gamma^{2}\log(z)(E-1)-4E/(1+\gamma)}{4E}}\right).$
When $r_{1}=r_{2}$,
$\left\|{\Delta_{zE}}\right\|_{\infty}=\frac{\left|r_{1}+r_{2}\right|}{2(1-\gamma)}\left|\beta_{E}^{z}\right|$
$\geq\frac{\left|r_{1}+r_{2}\right|}{2(1-\gamma)}\frac{E}{eT};$
otherwise,
$\|{\Delta_{zE}}\|_{\infty}$ $\geq\frac{\min\{\|(I-\bar{P})Q^{*}\|_{2},\|\bar{P}Q^{*}\|_{2}\}}{\sqrt{2}}\cdot\max\left\{|\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{z}\right\}$
$\geq\frac{1}{2}\min\left\{\left|r_{1}-r_{2}\right|,\frac{1}{1-\gamma}\left|r_{1}+r_{2}\right|\right\}\max\left\{|\alpha_{E}^{z}+\frac{1-\alpha_{E}^{z}}{1-\alpha_{E}}\kappa_{E}|,\beta_{E}^{z}\right\}$
$\geq\frac{1}{2}\min\left\{\left|r_{1}-r_{2}\right|,\frac{1}{1-\gamma}\left|r_{1}+r_{2}\right|\right\}\max\left\{\frac{1}{(1-\gamma)z}\left(\frac{\gamma^{2}\log(z)(E-1)-4E/(1+\gamma)}{4E}\right),\frac{1}{ez}\right\}$
$=\frac{1}{2}\min\left\{\left|r_{1}-r_{2}\right|,\frac{1}{1-\gamma}\left|r_{1}+r_{2}\right|\right\}\max\left\{\frac{E}{(1-\gamma)T}\left(\frac{\gamma^{2}\log(z)(E-1)-4E/(1+\gamma)}{4E}\right),\frac{E}{eT}\right\}.$
We can choose $\log(z)\geq\frac{4E(\gamma+2)}{(1+\gamma)\gamma^{2}(E-1)},E\geq 2$ so that $\left(\frac{\gamma^{2}\log(z)(E-1)-4E/(1+\gamma)}{4E}\right)\geq 1$. Then the first term inside the max operator is bigger. Then,
$\|\Delta_{zE}\|_{\infty}\geq\frac{1}{2}\min\left\{\left|r_{1}-r_{2}\right|,\frac{1}{1-\gamma}\left|r_{1}+r_{2}\right|\right\}\frac{E}{(1-\gamma)T}.$
@@ Appendix G Additional experiments
(a) Phase 2 stepsize $\lambda=\frac{1}{\sqrt{T}}$
(b) Phase 2 stepsize $\lambda_{t}=\frac{1}{\sqrt{t+1}}$
(c) Phase 2 stepsize $\lambda_{t}=\frac{c+1}{t+c}$, where $c=10$
(d) Phase 2 stepsize $\lambda_{t}=\frac{1}{t+1}$
(e) Phase 2 stepsize $\lambda_{t}=\frac{1}{(t+1)^{0.7}}$
@@
[…]
Figure 5: Homogeneous federated Q-learning with varying $E$.
G.2 Impacts of $E$ on time-decaying stepsize
Figure 6: Using time-decaying stepsize $\lambda_{t}=\frac{1}{\sqrt{t+1}}$, the overall convergence becomes worse as $E$ increases
Figure 7: Final error versus T. It is clear that when choosing $\lambda=\frac{1}{\sqrt{T}}$, the final error decays as $T$ increases.
G.3 Different target error levels.